OpenStudy (anonymous):

The equation and graph of a polynomial are shown below. The graph reaches its minimum when the value of x is 4. What is the y-value of this minimum? y=x^2-8x+12

OpenStudy (anonymous):

y = x^2-8x+12 put x=4 y= 4^2 - 8(4) +12 y= 16 - 32 +12 hence, y = -4